Antonio Banderos & Scarves makes headwear that is very popular in the fall-winter season. Units sold are anticipated as follows:

Monthly Unit Sales
October 1,100
November 2,100
December 4,200
January 3,200
10,600 Total units sold

If seasonal production is used, it is assumed that inventory will directly match sales for each month and there will be no inventory buildup.

However, Antonio decides to go with level production to avoid being out of merchandise. He will produce the 10,600 items over four months at a level of 2,650 per month.


a. What is the ending inventory at the end of each month? Compare the unit sales to the units produced and keep a running total.


b. If the inventory costs $2 per unit and will be financed at the bank at a cost of 12 percent, what is the monthly financing cost and the total for the four months? (Use 1 percent as the monthly rate.)
Answer #1
October November December January
i) Opening Inventory 0 1550 2100 550
ii) Units Produced 2650 2650 2650 2650
iii) Units Sold 1100 2100 4200 3200
iv) Closing Inventory (i+ii-iii) 1550 2100 550 0

b) Monthly Financing Cost = (Closing Inventory of that month) * (Cost per unit) * 1%

October = 1550 * 2 * 1% = $ 31

November = 2100 * 2 * 1% = $ 42

December = 550 * 2 * 1% = $ 11

January = 0 * 2 * 1% = 0

Total Financing Cost = $ (31+42+11+0) =$ 84